considering the two lines
y=2x-1 and y= -x +3
how do decide if the two lines intersect? and if they do what is the point of intersection? please show working. Thanks

Answer by user_dude2008(1862)

y=2x-1

-x+3=2x-1


3+1=2x+x

4=3x

x=4/3

y=2x-1=2(4/3)-1=8/3-1=5/3

Answers: x=4/3, y=5/3